### v1

@@ Abstract
Local search is a powerful heuristic in optimization and computer science, the complexity of which has been studied in the white box and black box models. In the black box model, we are given a graph $G=(V,E)$ and oracle access to a function $f:V\to\mathbb{R}$. The local search problem is to find a vertex $v$ that is a local minimum, i.e. with $f(v)\leq f(u)$ for all $(u,v)\in E$, using as few queries to the oracle as possible. We show that if a graph $G$ admits a lazy, irreducible, and reversible Markov chain with stationary distribution $\pi$, then the randomized query complexity of local search on $G$ is $\Omega\left(\frac{\sqrt{n}}{t_{mix}\cdot\exp(3\sigma)}\right)$, where $t_{mix}$ is the mixing time of the chain and $\sigma=\max_{u,v\in V(G)}\frac{\pi(v)}{\pi(u)}.$ This theorem formally establishes a connection between the query complexity of local search and the mixing time of the fastest mixing Markov chain for the given graph. We also get several corollaries that lower bound the complexity as a function of the spectral gap, one of which slightly improves a prior result from [BCR24].
@@ 1 Introduction
Local search stands as a robust heuristic within optimization and computer science, analyzed through both white box and black box frameworks. In the black box model, we are given a graph $G=(V,E)$ alongside oracle access to a function $f:V\to R$. The objective is to identify a vertex $v$ that represents a local minimum, meaning $f(v)\leq f(u)$ for every edge $(u,v)$, while minimizing the number of vertices queried. Obtaining lower bounds for the complexity of local search has a rich history of analysis via random walks. The first pioneering work on the subject was [Ald83], which did careful tailored analysis of the hitting time of random walks on the Boolean hypercube to obtain lower bounds for local search. Another breakthrough was obtained by [Aar06], which designed a combinatorial method of obtaining lower bounds for local search inspired by the relational adversary method from quantum computing. This approach enabled obtaining sharper lower bounds for the Boolean hypercube and $d$-dimensional grid, and was successfully used in many later works. In this paper we consider the high level question: How does the geometry of the graph affect the complexity of local search? While the query complexity is comprehensively understood for neighbourhood structures such as the $d$-dimensional grid and the Boolean hypercube, knowledge remains limited for more general neighbourhood structures. Nevertheless, the spatial structure in optimization settings typically extends to more complex graphs. For instance, in scenarios such as low rank matrix estimation with data compromised by adversarial attacks, the function is defined on a Riemannian manifold rather than a traditional Euclidean space [Bon13]; thus the discretization of an optimization search space may not necessarily always correspond to some $d$-dimensional grid. For a more extensive survey on stochastic gradient descent on Riemannian manifolds, see, e.g., [Ama13]. This motivates studying local search not only on hypercubes and grids, but also on broader classes of graphs. Inspired by the observation that many lower bounds for local search are based on various types of random walks, we consider general random walks for the graph at hand and obtain lower bounds as a function of their mixing time. Our analysis uses a variant of the classical relational adversary method from [BCR24] and our main result is generic in two ways: the graph is arbitrary and the random walk evolves according to a Markov chain, which we only require to be lazy, irreducible, and reversible. This allows us to formally connect the query complexity of local search and the mixing time of the fastest mixing Markov chain for the given graph, which is a classical problem analyzed starting with [BDX04], with recent results in [OTZ21]. As a corollary, we also get a lower bound in terms of the spectral gap of the transition matrix of the chain.
@@ 2.1 Model
Let $G=(V,E)$ be a connected undirected graph and $f:V\to\mathbb{R}$ a function defined on the vertices. A vertex $v\in V$ is a local minimum if $f(v)\leq f(u)$ for all $\{u,v\}\in E$. We will write $V=[n]=\{1,\ldots,n\}$. Given as input a graph $G$ and oracle access to function $f$, the local search problem is to find a local minimum of $f$ on $G$ using as few queries as possible. Each query is of the form: “Given a vertex $v$, what is $f(v)$?”.
@@
The deterministic query complexity of a task is the total number of queries necessary and sufficient for a correct deterministic algorithm to find a solution. The randomized query complexity is the expected number of queries required to find a solution with probability at least $9/10$ for each input, where the expectation is taken over the coin tosses of the protocol.
@@ 2.2 Our Contributions
[…]
The best lower bound given by 1 is attained by considering the Markov chain with the fastest possible mixing time for $G$; see [BDX04] for a classical reference on this problem. Indeed, for many classes of graphs there can significant gaps between the mixing time of the fastest mixing Markov chain and that of more obvious choices of Markov chains (such as the max-degree walk or the Metropolis-Hastings chain), including the barbell graph, edge-transitive graphs, and distance transitive graphs [BDPX09]. For example, when the stationary distribution is set to uniform on the barbell graph666two cliques of $n/2$ vertices connected by a single edge, the max-degree random walk has mixing time $\Theta(n^{3})$ while the fastest mixing walk mixes in only $\Theta(n^{2})$ steps (see corollaries $5.2$ and $5.3$ in [BDPX05]).
@@ 2.3 High Level Approach
The high level approach is as follows. We consider a set of value functions $f:V\to\mathbbm{R}$ induced by walks from a fixed starting vertex. We define $f$ such that the value at any vertex off the walk is the distance to the starting vertex of the walk, and the value at vertices on the walk is decreasing along the walk. This ensures that there is only one local minimum, namely the end of the walk. This type of construction is classical [Ald83, Aar06]. We denote the space of such functions $\mathcal{X}$. We then choose a similarity measure $r:\mathcal{X}\times\mathcal{X}\to\mathbbm{R}_{\geq 0}$ (also called the “relation”) between any two functions. Semantically, the relation measures the difficulty of distinguishing two possible input functions from each other; thus it will be useful that functions induced by walks with a long shared prefix are defined as more related. Given $r$, we invoke the relational adversary variant (Lemma 1 from [BCR24]), which implicitly defines a distribution over inputs based on $r$, and outputs a lower bound on the randomized query complexity. We choose $r$ carefully such that the distribution over walks is that of an arbitrary Markov chain. The innovation of our methodology over [BCR24] lies in the construction; where their construction is based on low-congestion paths, ours takes the more natural and general approach of using arbitrary Markov chains, including the usual lazy random walks.
@@
A key step in analyzing the formula given by the relational adversary is to bound the probability that, when fixing one staircase $\mathbf{x}$ and sampling a second staircase $\mathbf{y}$ conditioned to share a uniformly randomly long initial segment with $\mathbf{x}$, no vertex $v$ is too likely to lie on the “tail” of $\mathbf{y}$ (i.e. the portion after the initial segment shared with $\mathbf{x}$). The difference between our setting and previous random walk based methods lies in this analysis. We analyze separately the portions of the tail before and after it mixes. After $\mathbf{y}$ mixes, it is of course close to uniformly random. Before $\mathbf{y}$ mixes, the randomness of the point on $\mathbf{x}$ at which $\mathbf{y}$ diverges is enough to keep $\mathbf{y}$ close enough to uniformly random. See Figure 1 for a visual depiction. This analysis is very generic, parametrized only by the mixing time of the walk used. This allows 1 to give results for walks from arbitrary Markov chains with no additional analysis (other than estimating the mixing time). The natural mixing properties of the lazy random walk on expanders then allow us to derive strong lower bounds for such graphs.
@@ 2.4 Related Work
The query complexity of local search was first studied experimentally by [Tov81]. The first breakthrough in the theoretical analysis of local search was obtained by [Ald83]. Aldous stated the algorithm based on steepest descent with a warm start and showed the first nontrivial lower bound of $\Omega(2^{n/2-o(n)})$ on the query complexity for the Boolean hypercube $\{0,1\}^{n}$. The lower bound construction from [Ald83] uses Yao’s lemma and describes a hard distribution, such that if a deterministic algorithm receives a random function according to this distribution, the expected number of queries issued until finding a local minimum will be large. The random function is obtained as follows:
@@
The function $f$ defined this way has a unique local minimum at $v_{0}$. By a very delicate analysis of this distribution, [Ald83] showed a lower bound of $\Omega(2^{n/2-o(n)})$ on the hypercube $\{0,1\}^{n}$. This almost matches the query complexity of steepest descent with a warm start, which was also analyzed in [Ald83] and shown to take ${O}(\sqrt{n}\cdot 2^{n/2})$ queries in expectation on the hypercube. The steepest descent with a warm start algorithm applies to generic graphs too, resulting in ${O}(\sqrt{n\cdot d_{max}})$ queries overall for any graph with maximum degree $d_{max}$. Aldous’ lower bound for the hypercube was later improved by [Aar06] to $\Omega(2^{n/2}/n^{2})$ via a relational adversary method inspired from quantum computing. [Zha09] further improved this lower bound to $\Theta(2^{n/2}\cdot\sqrt{n})$ via a “clock”-based random walk construction, which avoids self-intersections. Meanwhile, [LTT89] developed a deterministic divide-and-conquer approach to solving local search that is theoretically optimal over all graphs for deterministic algorithms. On the hypercube, their method yields a lower bound of $\Omega(2^{n}/\sqrt{n})$ and an upper bound of ${O}(2^{n}\log(n)/\sqrt{n})$. For the $d$-dimensional grid $[n]^{d}$, [Aar06] used the relational adversary method there to show a randomized lower bound of $\Omega(n^{d/2-1}/\log n)$ for every constant $d\geq 3$. [Zha09] proved a randomized lower bound of $\Omega(n^{d/2})$ for every constant $d\geq 4$; this is tight as shown by Aldous’ generic upper bound. Zhang also showed improved bounds of $\Omega(n^{2/3})$ and $\Omega(n^{3/2}/\sqrt{\log n})$ for $d=2$ and $d=3$ respectively, as well as some quantum results. The work of [SY09] closed further gaps in the quantum setting as well as the randomized $d=2$ case. The problem of local search on the grid was also studied under the context of multiple limited rounds of adaptive interactions by [BL22]. [BDN19] studied the communication complexity of local search; this captures distributed settings, where data is stored on different computers.
@@
More general results are few and far between. On many graphs, the simple bound from [Ald83] of $\Omega(d_{max})$ queries is the best known lower bound: hiding the local minimum in one of the $d_{max}$ leaves of a star subgraph requires checking about half the leaves in expectation to find it. [SS04] gave a quantum lower bound of $\Omega\left(\sqrt[8]{\frac{s}{d_{max}}}/\log(n)\right)$, where $s$ is the separation number of the graph. This implies the same lower bound in a randomized context, using the spectral method. Meanwhile, the best known upper bound is ${O}((s+d_{max})\cdot\log n)$ due to [SS04], which was obtained via a refinement of the divide-and-conquer procedure of [LTT89]. [DR10] studied Cayley and vertex transitive graphs and gave lower bounds for local search as a function of the number of vertices and the diameter of the graph. [Ver06] obtained upper bounds as a function of the genus of the graph. [BCR24] improved on the relational adversary method of [Aar06] and applied it to get lower bounds for local search as a function of graph congestion, with corollaries for expansion and spectral gap. We apply their methodology to an alternate construction to achieve our results, including an improvement on their spectral gap result.
@@
Local search is strongly related to the problem of local optimization where the goal is to find a stationary point of a function on $\mathbb{R}^{d}$. A standard method for solving local optimization problems is through gradient methods, which find $\epsilon$-stationary points. To show lower bounds for finding stationary points, one can similarly define a function that selects a walk in the underlying space and hides a stationary point at the end of the walk. Handling (potential) smoothness of the function and ensuring the stationary point is unique represent additional challenges. For examples of works on algorithms and complexity of computing approximate stationary points, see, e.g., [Vav93, ${}^{+}$20, CDHS20, CDHS21, BM20, DS20]).
@@
The computational complexity of local search is captured by the class PLS, defined by [JPY88] to model the difficulty of finding locally optimal solutions to optimization problems. A related class is PPAD, introduced in [Pap94] to study the computational complexity of finding a Brouwer fixed-point. Both PLS and PPAD are subsets of the class TFNP. The class PPAD contains many natural problems that are computationally equivalent to finding a Brouwer fixed point ([CD09]), such as finding an approximate Nash equilibrium in a multiplayer or two-player game ([DGP09, CDT09]), an Arrow-Debreu equilibrium in a market ([VY11, CPY17]), and a local min-max point ([DSZ21]). The query complexity of computing an $\varepsilon$-approximate Brouwer fixed point was studied in a series of papers starting with [HPV89], which introduced a construction where the function is induced by a hidden walk. This was later improved by [CD05] and [CT07]. Recently, [FGHS22] showed that the class CLS, introduced by [DP11] to capture continuous local search, is equal to PPAD $\cap$ PLS.

### v3

@@ Abstract
Local search is a powerful heuristic in optimization and computer science, the complexity of which has been studied in the white box and black box models. In the black box model, we are given a graph $G=(V,E)$ and oracle access to a function $f:V\to\mathbb{R}$. The local search problem is to find a vertex $v$ that is a local minimum, i.e. with $f(v)\leq f(u)$ for all $(u,v)\in E$, using as few queries to the oracle as possible.
We show that if a graph $G$ admits a lazy, irreducible, and reversible Markov chain with stationary distribution $\pi$, then the randomized query complexity of local search on $G$ is $\Omega\left(\frac{\sqrt{n}}{t_{mix}\cdot\exp(3\sigma)}\right)$, where $t_{mix}$ is the mixing time of the chain and $\sigma=\max_{u,v\in V(G)}\frac{\pi(v)}{\pi(u)}.$ This theorem formally establishes a connection between the query complexity of local search and the mixing time of the fastest mixing Markov chain for the given graph. We also get several corollaries that lower bound the complexity as a function of the spectral gap, one of which slightly improves a prior result from [BCR24].
@@ 1 Introduction
Local search stands as a robust heuristic within optimization and computer science, analyzed through both white box and black box frameworks. In the black box model, we are given a graph $G=(V,E)$ alongside oracle access to a function $f:V\to R$. The objective is to identify a vertex $v$ that represents a local minimum, meaning $f(v)\leq f(u)$ for every edge $(u,v)$, while minimizing the number of vertices queried.
Obtaining lower bounds for the complexity of local search has a rich history of analysis via random walks. The first pioneering work on the subject was [Ald83], which did careful tailored analysis of the hitting time of random walks on the Boolean hypercube to obtain lower bounds for local search. Another breakthrough was obtained by [Aar06], which designed a combinatorial method of obtaining lower bounds for local search inspired by the relational adversary method from quantum computing. This approach enabled obtaining sharper lower bounds for the Boolean hypercube and $d$-dimensional grid, and was successfully used in many later works.
In this paper we consider the high level question: How does the geometry of the graph affect the complexity of local search? While the query complexity is comprehensively understood for neighbourhood structures such as the $d$-dimensional grid and the Boolean hypercube, knowledge remains limited for more general neighbourhood structures.
Nevertheless, the spatial structure in optimization settings typically extends to more complex graphs. For instance, in scenarios such as low rank matrix estimation with data compromised by adversarial attacks, the function is defined on a Riemannian manifold rather than a traditional Euclidean space [Bon13]; thus the discretization of an optimization search space may not necessarily always correspond to some $d$-dimensional grid. For a more extensive survey on stochastic gradient descent on Riemannian manifolds, see, e.g., [Ama13]. This motivates studying local search not only on hypercubes and grids, but also on broader classes of graphs.
Inspired by the observation that many lower bounds for local search are based on various types of random walks, we consider general random walks for the graph at hand and obtain lower bounds as a function of their mixing time. Our analysis uses a variant of the classical relational adversary method from [BCR24] and our main result is generic in two ways: the graph is arbitrary and the random walk evolves according to a Markov chain, which we only require to be lazy, irreducible, and reversible. This allows us to formally connect the query complexity of local search and the mixing time of the fastest mixing Markov chain for the given graph, which is a classical problem analyzed starting with [BDX04], with recent results in [OTZ21]. As a corollary, we also get a lower bound in terms of the spectral gap of the transition matrix of the chain.
@@ 2.1 Model
Let $G=(V,E)$ be a connected undirected graph and $f:V\to\mathbb{R}$ a function defined on the vertices. A vertex $v\in V$ is a local minimum if $f(v)\leq f(u)$ for all $\{u,v\}\in E$. We will write $V=[n]=\{1,\ldots,n\}$.
Given as input a graph $G$ and oracle access to function $f$, the local search problem is to find a local minimum of $f$ on $G$ using as few queries as possible. Each query is of the form: “Given a vertex $v$, what is $f(v)$?”.
@@
The deterministic query complexity of a task is the total number of queries necessary and sufficient for a correct deterministic algorithm to return a solution. The randomized query complexity is the expected number of queries required to return a solution with probability at least $9/10$ for each input, where the expectation is taken over the coin tosses of the protocol.
@@ 2.2 Our Contributions
$\Omega\left(\frac{\sqrt{n}}{t_{mix}\left(\frac{\sigma}{2n}\right)\cdot\exp(3\sigma)}\right),\qquad\;\mbox{where}\;\;\sigma=\max_{u,v\in V}\frac{\pi(v)}{\pi(u)}\,.$
The best lower bound given by 1 is attained by considering the Markov chain with the fastest possible mixing time for $G$; see [BDX04] for a classical reference on this problem.
Indeed, for many classes of graphs there can be significant gaps between the mixing time of the fastest mixing Markov chain and that of more obvious choices of Markov chains (such as the max-degree walk or the Metropolis-Hastings chain), including the barbell graph, edge-transitive graphs, and distance transitive graphs [BDPX09]. For example, when the stationary distribution is set to uniform on the barbell graph [^6] The barbell graph consists of two cliques of $n/2$ vertices each, connected by a single edge., the max-degree random walk has mixing time $\Theta(n^{3})$ while the fastest mixing walk mixes in only $\Theta(n^{2})$ steps (see corollaries $5.2$ and $5.3$ in [BDPX05]).
Do not be alarmed by the exponential dependence on $\sigma$. There are natural choices for $\mathcal{P}$ for which $\sigma=1$, such as the max-degree walk. Writing 1 in this way, rather than enforcing $\sigma=1$, gives a little flexibility for small (ideally constant) $\sigma$.
@@ 2.3 High Level Approach
The high level approach is as follows. We consider a set of value functions $f:V\to\mathbbm{R}$ induced by walks from a fixed starting vertex. We define $f$ such that the value at any vertex off the walk is the distance to the starting vertex of the walk, and the value at vertices on the walk is decreasing along the walk. This ensures that there is only one local minimum, namely the end of the walk. This type of construction is classical [Ald83, Aar06]. We denote the space of such functions $\mathcal{X}$.
We then choose a similarity measure $r:\mathcal{X}\times\mathcal{X}\to\mathbbm{R}_{\geq 0}$ (also called the “relation”) between any two functions. Semantically, the relation measures the difficulty of distinguishing two possible input functions from each other; thus it will be useful that functions induced by walks with a long shared prefix are defined as more related. Given $r$, we invoke the relational adversary variant (Lemma 1 from [BCR24]), which implicitly defines a distribution over inputs based on $r$, and outputs a lower bound on the randomized query complexity. We choose $r$ carefully such that the distribution over walks is that of an arbitrary Markov chain.
The innovation of our methodology over [BCR24] lies in the construction; where their construction is based on low-congestion paths, ours takes the more natural and general approach of using arbitrary Markov chains, including the usual lazy random walks.
@@
A key step in analyzing the formula given by the relational adversary is the following. When fixing a staircase $\mathbf{x}$ and sampling a second staircase $\mathbf{y}$ conditioned to share an initial prefix of random length with $\mathbf{x}$, the proof has to show that no vertex $v$ is too likely to lie on the “tail” of $\mathbf{y}$ (i.e. the portion after the initial segment shared with $\mathbf{x}$). The difference between our setting and previous random walk based methods lies in this analysis.
We analyze separately the portions of the tail before and after it mixes. After $\mathbf{y}$ mixes, it is close to being distributed according to the stationary distribution $\pi$. Before $\mathbf{y}$ mixes, the randomness of the point on $\mathbf{x}$ at which $\mathbf{y}$ diverges suffices to keep $\mathbf{y}$ close enough to being distributed according to $\pi$. A visual depiction is shown in Figure 1.
This analysis is very generic, parameterized only by the stationary distribution of the walk used. This allows 1 to give results for walks from arbitrary Markov chains with no additional analysis other than estimating the mixing time. The natural mixing properties of the lazy random walk on expanders then allow us to derive strong lower bounds for such graphs.
@@ 2.4 Related Work
The query complexity of local search was first studied experimentally by [Tov81]. The first breakthrough in the theoretical analysis of local search was obtained by [Ald83]. Aldous stated the algorithm based on steepest descent with a warm start and showed the first nontrivial lower bound of $\Omega(2^{n/2-o(n)})$ on the query complexity for the Boolean hypercube $\{0,1\}^{n}$.
The lower bound construction from [Ald83] uses Yao’s lemma and describes a hard distribution, such that if a deterministic algorithm receives a random function according to this distribution, the expected number of queries issued until finding a local minimum will be large. The random function is obtained as follows:
@@
The function $f$ defined this way has a unique local minimum at $v_{0}$. By a very delicate analysis of this distribution, [Ald83] showed a lower bound of $\Omega(2^{n/2-o(n)})$ on the hypercube $\{0,1\}^{n}$. This almost matches the query complexity of steepest descent with a warm start, which was also analyzed in [Ald83] and shown to take ${O}(\sqrt{n}\cdot 2^{n/2})$ queries in expectation on the hypercube. The steepest descent with a warm start algorithm applies to generic graphs too, resulting in ${O}(\sqrt{n\cdot d_{max}})$ queries overall for any graph with maximum degree $d_{max}$.
Aldous’ lower bound for the hypercube was later improved by [Aar06] to $\Omega(2^{n/2}/n^{2})$ via a relational adversary method inspired from quantum computing. [Zha09] further improved this lower bound to $\Theta(2^{n/2}\cdot\sqrt{n})$ via a “clock”-based random walk construction, which avoids self-intersections. Meanwhile, [LTT89] developed a deterministic divide-and-conquer approach to solving local search that is theoretically optimal over all graphs for deterministic algorithms. On the hypercube, their method yields a lower bound of $\Omega(2^{n}/\sqrt{n})$ and an upper bound of ${O}(2^{n}\log(n)/\sqrt{n})$.
For the $d$-dimensional grid $[n]^{d}$, [Aar06] used the relational adversary method there to show a randomized lower bound of $\Omega(n^{d/2-1}/\log n)$ for every constant $d\geq 3$. The relational adversary method that we use is inspired by Aaronson’s classical relational adversary. [Zha09] proved a randomized lower bound of $\Omega(n^{d/2})$ for every constant $d\geq 4$; this is tight as shown by Aldous’ generic upper bound. Zhang also showed improved bounds of $\Omega(n^{2/3})$ and $\Omega(n^{3/2}/\sqrt{\log n})$ for $d=2$ and $d=3$ respectively, as well as some quantum results. The work of [SY09] closed further gaps in the quantum setting as well as the randomized $d=2$ case.
The problem of local search on the grid was also studied under the context of multiple limited rounds of adaptive interactions by [BL22]. [BDN19] studied the communication complexity of local search; this captures distributed settings, where data is stored on different computers.
@@
More general results are few and far between. On many graphs, the simple bound from [Ald83] of $\Omega(d_{max})$ queries is the best known lower bound: hiding the local minimum in one of the $d_{max}$ leaves of a star subgraph requires checking about half the leaves in expectation to find it. [SS04] gave a quantum lower bound of $\Omega\left(\sqrt[8]{\frac{s}{d_{max}}}/\log(n)\right)$, where $s$ is the separation number of the graph. This implies the same lower bound in a randomized context, using the spectral method. Meanwhile, the best known upper bound is ${O}((s+d_{max})\cdot\log n)$ due to [SS04], which was obtained via a refinement of the divide-and-conquer procedure of [LTT89].
[DR10] studied Cayley and vertex transitive graphs and gave lower bounds for local search as a function of the number of vertices and the diameter of the graph. [Ver06] obtained upper bounds as a function of the genus of the graph. [BCR24] improved on the relational adversary method of [Aar06] and applied it to get lower bounds for local search as a function of graph congestion, with corollaries for expansion and spectral gap. We apply their methodology to an alternate construction to achieve our results, including an improvement on their spectral gap result.
@@
Local search is strongly related to the problem of local optimization where the goal is to find a stationary point of a function on $\mathbb{R}^{d}$. A standard method for solving local optimization problems is through gradient methods, which find $\epsilon$-stationary points. To show lower bounds for finding stationary points, one can similarly define a function that selects a walk in the underlying space and hides a stationary point at the end of the walk. Handling (potential) smoothness of the function and ensuring the stationary point is unique represent additional challenges. For examples of works on algorithms and complexity of computing approximate stationary points, see, e.g., [Vav93, ZLJ+20, CDHS20, CDHS21, BM20, DS20]).
@@
The computational complexity of local search is captured by the class PLS, defined by [JPY88] to model the difficulty of finding locally optimal solutions to optimization problems. A related class is PPAD, introduced in [Pap94] to study the computational complexity of finding a Brouwer fixed-point. Both PLS and PPAD are subsets of the class TFNP.
The class PPAD contains many natural problems that are computationally equivalent to finding a Brouwer fixed point ([CD09]), such as finding an approximate Nash equilibrium in a multiplayer or two-player game ([DGP09, CDT09]), an Arrow-Debreu equilibrium in a market ([VY11, CPY17]), and a local min-max point ([DSZ21]). The query complexity of computing an $\varepsilon$-approximate Brouwer fixed point was studied in a series of papers starting with [HPV89], which introduced a construction where the function is induced by a hidden walk. This was later improved by [CD05] and [CT07]. Recently, [FGHS22] showed that the class CLS, introduced by [DP11] to capture continuous local search, is equal to PPAD $\cap$ PLS.